a. A certain tractor is said to be capable of pulling with a steady force of 18kN while moving at a velocity of 8.0m/s. How much power in kilowatts and in horsepower is the tractor developing under these conditions?



1
Expert's answer
2021-05-21T17:06:06-0400

"Power = forces \times velocity \\
P = 18000 \times 8 \\
P = 144000 = 144 \;kW
1 \;kw = 1.3 \; horsepower \\
144 \;kw = 1.3 \times 144 = 187.2 \;hp"